OpenStudy (anonymous):

solve x+>-3/11 write your solution using set builder notation. please show all work. the> is underlined

OpenStudy (anonymous):

Hi! What is confusing you about this problem?

OpenStudy (paxpolaris):

\[x \ge- {3\over11}\] Is this what the problem looks like: why is there a + after x?

OpenStudy (anonymous):

thats how the problem is written

OpenStudy (anonymous):

may be it's a mistake

OpenStudy (anonymous):

you have to solve problem first than write it in set builder notation and the problem is x+2=-3/11

OpenStudy (paxpolaris):

you can type ‘>=’ for the “greater than or equal to” sign (≥)

OpenStudy (paxpolaris):

\[x+2 \ge -{3 \over 11}\] So, \[x \ge -{3 \over 11}-2\] \[x \ge -2{3 \over 11}\] or \[x \ge -{25 \over 11}\]

OpenStudy (paxpolaris):

IF there is no other condition, like x is an integer, or x is a natural number: \[\left\{ x: x \in \mathbb{R}, \&\ x \ge -2{3\over11}\right\}\]

OpenStudy (anonymous):

where did you get 25 from

OpenStudy (paxpolaris):

\[2{3\over11} = {2\times 11\over 11} +{3\over11}={22\over 11} +{3\over11}={22+3\over11}={25\over11}\]

OpenStudy (paxpolaris):

\[-2{3\over11} =-{25\over11}\] You could write the answer in set builder form just as : \[\left\{ x :x ≥-2 {3\over11} \right\} \]